i need help solving this problem.... please and thank you... jim runs a food cart and during a busy out door festival he sold $7343.75 worth of food. he sells hot dogs for $2.95 and steak sandwiches for $9.95. if he sold a total of 985 items that day,how many of each item did he sell?
Answer by stanbon(75887)

jim runs a food cart and during a busy out door festival he sold $7343.75 worth of food. he sells hot dogs for $2.95 and steak sandwiches for $9.95. if he sold a total of 985 items that day,how many of each item did he sell?
--------------------------------------------
Quantity Equation: d + s = 985
Value Equation::295d + 995s = 734375 cents
--------------------------------------------------
Multiply thru the 1st equation by 295 to set up the system for elimination:
295d + 295s = 295*985
295d + 995s = 734375 cents
----------------------
Subtract the 1st from the 2nd so you can solve for "s":
700s = 443800
s = 634 (# of steack sandwiches sold)
----------------
Substitute into d+s = 985 to solve for "d":
d + 634 = 985
d = 351 (# of hot dogs sold)
=================================
Cheers,
Stan H.